The length of rectangle is 5m greater than the width. The perimeter is 150m more then twice the width. Find the width and length of the rectangle. Please explain and show how to do it thanks:)
Answer by checkley77(12844)

L=W+5
P=2W+150
P=2W+2L
2L=150 REPL;ACE L WITH (W+5) & SOLVE FOR W.
2(W+5)=150
2W+10=150
2W=150-10
2W=140
W=140/2
W=70 ANS. FOR THE WIDTH.
L=70+5=75 ANS FOR THE LENGTH.
P=2*70+2*75
P=140+150
P=290 ANS. FOR THE PERIMETER.